OpenStudy (crashonce):

Find the A.P. in which S (10) = 465 and 9 S(3) = 4 S(6) Show working

OpenStudy (crashonce):

@ikram002p @ganeshie8

OpenStudy (perl):

AP stands for arithmetic progression?

OpenStudy (crashonce):

yes

OpenStudy (perl):

here is a formula for sum of AP

OpenStudy (perl):

that should get you started

OpenStudy (crashonce):

but i dont know how to go from there

OpenStudy (perl):

we use the formula S(n) = n*(a1 + an) / 2

OpenStudy (perl):

so we are given S(10) = 465

OpenStudy (perl):

S(10) = 465 S(10) = 10 * (a1 + an) /2 so 465 = 10*(a1+an)/2

OpenStudy (perl):

the next thing it says 9 S(3) = 4 S(6) , im assuming that means multiplication

OpenStudy (perl):

9*S(3) = 4* S(6)

OpenStudy (perl):

9 * 3 * (a1 + an)/2 = 4 * 6 (a1 +an)/2

OpenStudy (perl):

so far we have two equations 465 = 10*(a1+an)/2 9 * 3 * (a1 + an)/2 = 4 * 6 (a1 +an)/2 we can solve a system of two linear equations in 2 unknowns

OpenStudy (perl):

wolfram says there is no solution, so we might be doing something wrong http://www.wolframalpha.com/input/?i=solve+465+%3D+10*%28a%2Bb%29%2F2%2C+9+*+3+*+%28a+%2B+b%29%2F2+%3D+4+*+6+%28a+%2Bb%29%2F2+

OpenStudy (crashonce):

yea

OpenStudy (perl):

can you upload your worksheet or take a screenshot

OpenStudy (crashonce):

Questions correct cant take a screenshot